OpenStudy (anonymous):

How do I get (2x^2-1)/(x(x^2-1)) from... y'=(1/x)+(1/2)*(2x/(x^2-1))

OpenStudy (anonymous):

\[y'=\frac{ 1 }{ x }+ \frac{ 1 }{ 2 } (\frac{ 2x }{ x^2-1 })\]